Calculate the grams of O 6.64 g Li3PO4



1
Expert's answer
2021-09-24T02:23:19-0400

Since lithium orthophosphate contains 3 lithium ions, the amount of lithium ion substance will be equal to:

"n(Li3PO4) = \frac{6.64}{115.79} = 0.0573\ mole"

"n(Li^+) = n(Li3PO4) \times 3 = 0.172 \ mole"